solve: the width of a rectangle is 3m less than 2/3 the length, perimeter 114. solve for width and length.
Answer by rfer(16322)

2x+2(2/3x-3)=114
2x+4/3x-6=114
3 1/3x=120
x=120/(10/3)
invert & multiply
x=360/10
reduce
x=36
2/3x-3=21